OpenStudy (anonymous):

1.What's it means that "Find the curves that are perpendicular to the parabolas y = ax^2"? 2.how to prove that "if a line has slope m, a line perpendicular to it will have slope −1/m"?

OpenStudy (anonymous):

1. perpendicular means the tangent lines are orthogonal at intersection point. 2. We may use vectors. (1,m) dot (1,-1/m) =0.

OpenStudy (anonymous):

Tangent lines means each line of both curves.

OpenStudy (anonymous):

2. if a is the angle line 1 makes with x axis, then its slope m = tan (a). Now, let's suppose the angle with the x-axis of another line is b such that tan (b) = -1/m. Your question is how come a = b-pi/2? Remember two trig identities: tan(-x) = -tan (x) [identity 1], tan (pi/2 -x)= cot (x) = 1/tan(x) [identity 2] so, -tan (x-pi/2) = 1/tan(x). (I applied identity 1 on identity 2] or, tan (x-pi/2) = -1/tan(x) Now putting a for x, we get tan (a-pi/2) = -1/tan(a) = -1/m which we had assumed was tan (b) hence, a-pi/2 = b

OpenStudy (anonymous):

I believe that the first part of this question is an orthogonal trajectories problem.